An electron has a mass of 9.11×10^−31 kg. The charge on an electron is −1.60×10^−19 C .

An electron is released from rest in a field that has a magnitude of 100 V/m = 100 N/C. What is the magnitude of the electron's acceleration, in m/s/s? Please provide your answer both in scientific notation and as a decimal.
Answer #1

Solution) m = 9.11×10^(-31) kg

q = - 1.6×10^(-19) C

E = 100 N/C

a = ?

Force , F = E|q|

ma = E|q|

a = (E|q|)/(m)

a = (100×1.6×10^(-19))/(9.11×10^(-31))

a = 17.56×10^(12) m/s^2

Therefore , a = 17.6×10^(12) m/s^2